OpenStudy (anonymous):

how many permutations are there of 3 things taken from a set of 7?

OpenStudy (anonymous):

@amistre64

OpenStudy (nubeer):

if you are using a calculator.. you can just plug in the values 7P3

OpenStudy (amistre64):

the way i remember the is to start at 7 and pick 3 numbers to multiply: 7*6*5 = 7P3

OpenStudy (anonymous):

where does the 6 come from?

OpenStudy (amistre64):

7 ways to pick the first element 6 ways to pick the second element 5 ways to pick the third element

OpenStudy (amistre64):

if we made a tree diagram of this, it would be 7 branches with 6 branches each; 7*6 and each branch has 5 leaves: (7*6)*5 total outcomes

OpenStudy (amistre64):

for any given set of elements, there are 6 ways that they can be presented: abc acb bac bca cab cba, which means that for combinations, we have 6 like sets. (7P3)/3! would be the combinatorial calculation ... but thats not really what it asked for is it :)

OpenStudy (anonymous):

so confused. whats the simplest way to solve this?

OpenStudy (amistre64):

start at 7 and pick 3 numbers to multiply: 7*6*5

OpenStudy (anonymous):

so the answer is 210? i dont get how

OpenStudy (anonymous):

how do you pick the 3 numbers?

OpenStudy (amistre64):

i explained how ... which you say is confusing 6*5 = 30, 30*7 = 210

OpenStudy (amistre64):

start at 7 and go down

OpenStudy (amistre64):

spose you have a litter of 7 kittens; how many ways can you choose the first one to pick up? after picking up the first one, how many ways are left in which you can choose to pick up the next one? after youve picked 2 kittens, how many ways are left for you to be able to choose the last kitten?

OpenStudy (anonymous):

so if it was: how many permutations are there of 2 things taken from a set of 5? the answer would be: 5*4 =20 ?

OpenStudy (amistre64):

7 to choose from; then 6 to choose from, then 5 to choose from

OpenStudy (amistre64):

yes 5*4 is the permutations of 2 things from 5

OpenStudy (anonymous):

ohh thank youu

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

"If K objects out of a set of N elements are placed into a fixed order, there are ________ unique arrangements of those objects" ??

OpenStudy (amistre64):

start with n and start picking out k elements or rather, theres some formula you can use ....\[\frac{n!}{(n-k)!}\]

OpenStudy (anonymous):

huhh??

OpenStudy (kropot72):

Are the N elements all different?

OpenStudy (amistre64):

a set is properly defined by distinct elements ....

OpenStudy (amistre64):

a set that has N elements is by definition N different elements

OpenStudy (anonymous):

so the blank would be N, correct?

OpenStudy (amistre64):

no

OpenStudy (amistre64):

7 Pick 3 did not amount to 7

OpenStudy (amistre64):

\[\frac{7*6*5*4*3*2*1}{4*3*2*1}=\frac{7!}{(7-3)!}\]

OpenStudy (anonymous):

i dont get the bottom..

OpenStudy (amistre64):

k elements picked from a set of n elements is usually calculated as: n!/(n-k)!

OpenStudy (amistre64):

notice that in order to simplify this to 7*6*5, we have to divide out the 4*3*2*1

OpenStudy (amistre64):

in general; to simplify n! to account for nPk, we need to divide out (n-k)!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so....

OpenStudy (anonymous):

soo lost...